OpenStudy (anonymous):

In a graduating class with the same number of boys and girls, 1/8 of the girls and 5/6 of the boys are honor students. What part of the class consist of boys who are not honor students?

OpenStudy (anonymous):

8/8-1/8=7/8 7/8-5/6=42/48-40/48 =2/48 = 1/14